OpenStudy (anonymous):

Jeff stores coffee powder in a steel container with a capacity of 252.3 cubic inches. He uses a glass container to store sugar. The two containers are congruent cylinders. If the glass container has a height of 6.2 inches, what is its radius? Answer 1.8 inches 2.1 inches 3.1 inches 3.6 inches

hero (hero):

Hint: Use Volume of Cylinder formula

OpenStudy (anonymous):

V = pi*h*r^2 252.3 = 6.2pi *r^2 r = sqrt(252.3/6.2pi)